OpenStudy (anonymous):

What is the surface area of the square pyramid? 48 cm2 144 cm2 156 cm2 96 cm2

OpenStudy (anonymous):

pyramid

Directrix (directrix):

The surface area of a square-based pyramid is 1/2 times the perimeter of the base times the slant height of the pyramid PLUS the base area which is 6*6.

Directrix (directrix):

SA = 1/2 * (4*6)*5 + 6*6 SA = ? @Joshuam1

OpenStudy (anonymous):

96

OpenStudy (anonymous):

thanks :D 4 more

OpenStudy (anonymous):

same pyramid but different question What is the volume of the square pyramid? 48 cm3 144 cm3 156 cm3 96 cm3

Directrix (directrix):

The volume of the square pyramid is (1/3) B* h where B is the area of the square base and height is the height of the solid. V = (1/3) * 6*6*4 V = ?

OpenStudy (anonymous):

48

Directrix (directrix):

In cubic measure, yes, that is what I got.

OpenStudy (anonymous):

thanks 3 more

OpenStudy (anonymous):

i'm gonna make a new post for this one

Directrix (directrix):

New thread time.